OpenStudy (anonymous):

helpp... Betty has 10 more dimes than quarters. If she has $3.45, how many coins does she have?

OpenStudy (anonymous):

d=dimes q=quarters d=q+10 3.45 = 0.1d + 0.25q 3.45 = 0.1(q + 10) + 0.25q 3.45 = 0.1q + 1 + 0.25q 3.45 - 1 = 0.1q + 0.25q 2.45 = 0.35q 7=q d=q + 10 d= 7 + 10 d=17 7 quarters and 17 dimes

OpenStudy (anonymous):

thank u so much....